OpenStudy (anonymous):

The circulation of a newsletter decreased from 3,200 to 2,464. What was the percent of decrease in circulation? A. 129% B. 77% C. 2.3% D. 23%

sam (.sam.):

D. 23%

sam (.sam.):

3200-2464=736 (736/3200)*100=23%